OpenStudy (anonymous):

Which of the following equations has the given solution set? Solution set: Ø -2m + 5 = 2m + 5 -2m + 5 = -2m + 5 -2m + 5 = -2m - 5 2m + 5 = 2m + 5 Ermmm...WHAT? @Ashleyisakitty

OpenStudy (ashleyisakitty):

The answer, i believe, you are looking for is choice A. I plugged them all into my calculator and the only one to come up with a probably graph was A, and theta was only to do with graphs, so im assuming no other answer except A could work.

OpenStudy (anonymous):

Thank you ashley...ive never even seen that before and it was on a quiz...That is just lovely isnt it?

OpenStudy (ashleyisakitty):

Seeing things on quizzes ive never seen before is my absolute favorite ;D

OpenStudy (anonymous):

Its lovely when its on the end of course test too isnt it?

OpenStudy (ashleyisakitty):

the beeesssttt

OpenStudy (anonymous):

Luckily the eoct is on my computer and i managed to get into the wifi.

OpenStudy (anonymous):

Ashley...one more quick question... Which of the following would eliminate the variable on the left side of the given equation? 36 - 19w = -6w + 41 add 19w correct?

myininaya (myininaya):

empty set means there are no solutions the answer isn't choice A

OpenStudy (anonymous):

But myiniaya there isnt a option for no solutions...IM SCREWED D:

myininaya (myininaya):

Yes one of the equations will lead to a no solution answer.

myininaya (myininaya):

But it isn't the first one.

OpenStudy (anonymous):

so...which one...and please hurry...eoct time is running slim...

myininaya (myininaya):

-2m + 5 = 2m + 5 -2m + 5 = -2m + 5 -2m + 5 = -2m - 5 2m + 5 = 2m + 5 ----------------- Do you know about lines? Like you know parallel lines never intersect right? Parallel lines have the same slope but different y-intercept. Looking at both sides of the equation you have something like am+b (where a is the slope and b is the y-intercept) Do any of these have the same a on both sides but different b.

OpenStudy (anonymous):

>_> ermmm Ok...I understand that part but...they never taught us slope and stuff yet...Please assist...

myininaya (myininaya):

Ok. Then solve equation equation for m. If you end up with something that can't be true in your eyes, then there is no solution. Like we know 1 doesn't equal 0. So if you end up with a situation like this, then you know there no solution. If we end up with something that is always true like 2=2 or 0=0, then there is infinite amount of solutions. There is the other case where we will end up with just the one solution. m=something.

myininaya (myininaya):

So when you solve the first equation for m, what do we get?

myininaya (myininaya):

-2m + 5 = 2m + 5 Get your m's on one side and your constants on the other side.